# PlatoNeRF: 3D Reconstruction in Plato’s Cave via Single-View Two-Bounce Lidar

### 21 Dec 2023 v1

@@ 4.2 Results
Simulated Results. Depth metrics for both the training view and across 120 test views for both our method and the baseline methods are reported in Table 1, and Chamfer distance is reported in Table 2. We find that our method consistently outperforms both BF Lidar and $\text{S}^{3}$-NeRF across both sets of metrics. Qualitative results are shown in Figure 4. Our method is able to reconstruct the visible and occluded parts of the scene, providing accurate scale and absolute depth. Due to our use of an implicit representation, we achieve much smoother results than BF Lidar. However, because our method uses vanilla NeRF, there are small floaters visible in some results. In contrast, $\text{S}^{3}$-NeRF uses an SDF representation, reducing floaters, but resulting in overly smoothed results, e.g. the dragon’s head.
@@
Figure 5: Real-World Results. (a) Captured scene (stars are illumination spots), (b) BF Lidar result, (c) PlatoNeRF result. Our method achieves similar results as BF Lidar, with much fewer artifacts.
Real-World Results. We show real-world results in Figure 5 and compare with BF Lidar. PlatoNeRF method achieves competitive performance. While BF Lidar produces many artifacts, especially near edges and specular areas on the mannequin, PlatoNeRF produces far fewer, despite not modeling specular surfaces. In general, PlatoNeRF produces smoother depth, but small floaters are noticeable, especially in the nearby floor region, which is an area for future work.
@@ 4.3 Ablations
We ablate our method to understand how it is affected by (1) reduced spatial and temporal resolution, (2) ambient light, (3) background albedo, (4) non-planar surfaces, and (5) number of illum. points. (1) is ablated in comparison to BF Lidar to highlight the benefits of our method over related lidar work. (2) and (3) are ablated in comparison to $\text{S}^{3}$-NeRF to highlight the fundamental advantages of using lidar compared to RGB when measuring shadows. Finally, (4) and (5) are done only on PlatoNeRF. All ablations are done on the chair scene.
@@ 4.3.1 Spatial and Temporal Resolution
In comparison to research-grade lidars, lidars on consumer devices have lower spatial and temporal resolution. We highlight an advantage of PlatoNeRF over BF Lidar, which is better generalization to low-resolution regimes due to our implicit representation. Quantitative and qualitative results are shown in Table 3 and Figure 6 (rows 1–2), respectively.
@@ 4.3.2 Ambient Light and Low Albedo Backgrounds
In real-world settings, there may be high ambient light or low scene albedo, both of which make detection of shadows in RGB images more challenging. In contrast, lidar-based methods, such as PlatoNeRF, are fundamentally more robust to these low signal-to-noise (SNR) and signal-to-background (SBR) scenarios. While ambient light and low albedo impact both RGB and lidar-based approaches, lidar-based approaches can mitigate their impact with time- and wavelength-gating. Ambient noise is uniformly distributed over time. Time gating enables suppression of ambient noise by only considering timing bins containing the pulse signal. Wavelength gating enables suppression of ambient noise in wavelengths not measured by the SPAD sensor, whereas RGB sensors have broadband sensitivity to visible wavelengths. A detailed explanation of the gating principle is provided in the supplement. Empirical results for these ablations are reported in Table 4.
@@
Increasing the ambient light would increase the background and noise in the SBR or SNR term. We empirically validate that PlatoNeRF is able to handle ambient light in the scene, while $\text{S}^{3}$-NeRF depth error increases. For this experiment, we render the scene with an added area light to train both methods. The area light intensity is the same in both RGB and lidar rendering. We do not model saturation or pileup distortion [30] effects, since these are not significant in indoor environments. PlatoNeRF’s reconstruction under ambient light is shown in Figure 6 (row 3) and $\text{S}^{3}$-NeRF’s is in the supplement.
@@
Reducing the albedo would reduce the signal in the SNR and SBR term. However, unlike RGB sensors, scene albedo has a minor effect on lidar. To show the impact of albedo on PlatoNeRF and $\text{S}^{3}$-NeRF, we reduce the albedo of all background surfaces. While the shadow is still discernible to the human eye (see supplement), the lowered albedo causes $\text{S}^{3}$-NeRF’s depth error to increase, while PlatoNeRF is unaffected. $\text{S}^{3}$-NeRF produces accurate depth from the training view, but, in occluded regions, geometry is missing due to the weaker shadows.
@@ 4.3.3 Other Ablations
Lastly, we study the impact of both non-planar background geometry and number of illumination spots on PlatoNeRF reconstructions. The qualitative results for both are shown in Figure 6 (row 3). The non-planar scene consists of curved walls on either side. PlatoNeRF’s depth L1 is 7.75 cm and PSNR is 27.25 dB across test views, similar to previous experiments. For our illumination spot ablation, we reduce the number of illumination spots from 16 to 8, leading to L1 error of 9.12 cm and PSNR of 26.33 dB across test views, a small drop from when all sixteen views are used. Varying illumination spots are studied more in the supplement.
@@ 5 Conclusion
We believe this research is a promising direction as lidar becomes ubiquitous. Future directions enabled by PlatoNeRF include incorporating RGB and lidar with neural rendering, incorporating data priors that are commonly used for single-view 3D reconstruction into PlatoNeRF, and modeling more than two-bounces of light.
@@ Appendix A Time & Wavelength Gating in Lidar
The principle of time gating is illustrated in Figure 7. A measured lidar signal $i(t)$ can be decomposed into the pulse signal $s(t)$ and (roughly) constant ambient background noise $n(t)=N$. An RGB sensor would integrate over this timing information and measure
@@
where $T_{1}$ and $T_{2}$ determine the gated window in the transient signal and $W=T_{2}-T_{1}$ is the window size. Note that the numerator of Equation 11 is roughly the same as the SBR in Equation 13 because $s(t)\approx 0$ for $t<T_{1}$ and $t>T_{2}$, as shown in Figure 7(a). Therefore, time gating offers an SNR improvement of $\frac{T}{W}$ over techniques that leverage RGB or intensity signals. Note that the SBR enhancement is inversely proportional to the gated window. We do not account for Poisson noise effects, which, in practice, would introduce trade-offs in determining the window size. Empirical results are plotted in Figure 7(b)-(d) on the effects of time gating on enhancing contrast in shadow images.
@@
A similar idea can be applied to gate wavelengths. Most of the signal will be concentrated within a narrow spectral range, and all other intensities can be gated out with a narrow-band pass filter, as shown in Figure 8. This figure plots the emission spectra of an LED light [36] and the gating profile is determined by a $685$ nm PicoQuant pulsed laser [31].
@@ Appendix B Simulated Dataset Details
In this section, we describe the simulated datasets that we render and use to compare our method to past work in more detail. We render four simulated scenes, as described in the main text, with both a lidar and RGB camera in Mitsuba [12]. The lidar data is used to run PlatoNeRF and Bounce Flash (BF) Lidar [7] and the RGB data is used to run $\text{S}^{3}$-NeRF [46]. The same sixteen scene points are illuminated in both the lidar and RGB data. In the lidar data, the sixteen points are illuminated with a laser and, in the RGB data, point light sources are placed at each of the sixteen points. A camera to world transform from OpenGL (x right, y up, z back) to Mitsuba (x left, y up, and z forward) is used to train each method with this data. Ground truth depth for both the train view and 120 test views are provided. A subset of the test view frames are shown in the video results on the project page. We plan to release all data for use in future work.
@@ Appendix C Training Details
We will release all data, code, and model checkpoints, along with documentation, prior to camera ready to ensure our work is fully reproducible by others. Code is provided in the ”code” folder of the supplement. Simulated data is rendered in Mitsuba world coordinates and PlatoNeRF uses OpenGL camera coordinates.
@@
We train our model for 200k iterations. For the first 25k iterations, only the distance loss is applied, while both the distance and shadow losses are applied thereafter. We use a threshold of 15% on the shadow confidence map (computed as the maximum of the cross-correlation described in Section 3.3) when extracting ground truth shadow masks from the raw lidar measurements. This threshold is used across all experiments, except the ambient light experiment, where we further tune it.
@@ Appendix D Extended Ablations
In this section, we add further detail and discussion on the results of our ablation, quantitatively reported in the main text. In addition, we provide further ablation on the impact of the number of illumination points on PlatoNeRF (Table 5).
@@ D.1 Spatial- and Temporal-Resolution
Figure 10: Ambient Light and Low Albedo Background Ablation. We compare PlatoNeRF and $\text{S}^{3}$-NeRF when trained on a scene with either ambient light or a low albedo background. PlatoNeRF is robust to both, whereas the performance of $\text{S}^{3}$-NeRF degrades.
Qualitative results comparing PlatoNeRF and Bounce Flash (BF) Lidar under varying spatial- and temporal-resolutions are shown in Figure 9. This ablation is important because lidars on consumer devices are often constrained to much lower resolutions than research-grade lidars. Spatial resolution is varied by downsampling the number of pixels, while keeping the field of view of the lidar the same. As spatial resolution is reduced, geometry predicted by BF Lidar becomes sparser. The depth estimation of visible points in the scene remains accurate, but there is no interpolation between these points. The sparsity in visible depth information negatively impacts the shadow carving step of BF Lidar, leading to poor reconstruction of the chair in lower spatial resolution regimes. On the other hand, because PlatoNeRF is able to smoothly interpolate across missing pixels, the resulting reconstruction is significantly more accurate.
@@ D.2 Ambient Light
Qualitative results comparing PlatoNeRF and $\text{S}^{3}$-NeRF reconstructions under ambient light are shown in Figure 10 (top row). While $\text{S}^{3}$-NeRF is able to model small amounts of ambient light, it fails under realistic amounts of ambient light, in this case, from an added area light. On the other hand, PlatoNeRF is still able to accurately reconstruct the scene with the same ambient light added.
@@ D.3 Low-Albedo Backgrounds
Qualitative results comparing PlatoNeRF and $\text{S}^{3}$-NeRF reconstructions with a low albedo background are shown in Figure 10 (bottom row). $\text{S}^{3}$-NeRF is able to accurately reconstruct the visible portion of the scene, but is unable to recover occluded geometry due to worse contrast in the shadow (though it is still discernible to the human eye, as shown in Figure 10). On the other hand, PlatoNeRF is not significantly affected by scene albedo due to its use of a lidar rather than RGB sensor.
D.4 Number of Illumination Points
We further ablate the impact of reducing the number of illumination points used to train PlatoNeRF. In our main experiments, we use sixteen illumination points. We reduce that number to eight, four, and two and report the results in Table 5. Qualitative results are shown in Figure 11. The scene is reconstructed for each number of illumination points, however, as the number is reduced, quality also decreases, as there is less information about occluded regions. When there are only two illumination points, the occluded chair legs are not reconstructed. We note that while we study the number of illumination points, their location is also an important factor in reconstruction quality. As the number of illumination points is reduced, the location of the remaining illumination points becomes increasingly important, i.e. casting shadows with the most relevance and diversity. In these experiments, we randomly choose which illumination points to use.

### 05 Apr 2024 v2

@@ 4.2 Results
Simulated Results. Depth metrics for the training and 120 test views for our method and the baseline methods are reported in Tab. 1, and Chamfer distance is reported in Tab. 2. We find that our method consistently outperforms both BF Lidar and $\text{S}^{3}$-NeRF across both sets of metrics. Qualitative results are shown in Fig. 4. Our method is able to reconstruct the visible and occluded parts of the scene, providing accurate scale and absolute depth. Due to our use of an implicit representation, we achieve much smoother results than BF Lidar. However, because our method uses vanilla NeRF, there are small floaters visible in some results. In contrast, $\text{S}^{3}$-NeRF uses an SDF representation, reducing floaters, but resulting in overly smoothed results, e.g. the dragon’s head.
@@
Figure 5: Real-World Results. (a) Captured scene (stars are illumination spots), (b) BF Lidar result, (c) PlatoNeRF result. Our method yields similar results as BF Lidar, with much fewer artifacts/holes.
Real-World Results. We show real-world results in Fig. 5 and compare with BF Lidar. PlatoNeRF method achieves competitive performance. While BF Lidar produces many artifacts and holes, especially near edges and specular areas on the mannequin, PlatoNeRF produces far fewer, despite not modeling specular surfaces. In general, PlatoNeRF produces smoother depth, but small floaters are noticeable, especially in the nearby floor region, which is an area for future work.
@@ 4.3 Ablations
We ablate our method to understand how it is affected by (1) reduced spatial and temporal resolution, (2) ambient light, (3) background albedo, (4) non-planar surfaces, (5) number of illumination points, and (6) shadow mask threshold. (1) is ablated in comparison to BF Lidar to highlight the benefits of our method over related lidar work. (2) and (3) are ablated in comparison to $\text{S}^{3}$-NeRF to highlight the fundamental advantages of using lidar compared to RGB when measuring shadows. Finally, (4), (5), and (6) are done only on PlatoNeRF. All ablations are done on the chair scene.
@@ 4.3.1 Spatial and Temporal Resolution
In comparison to research-grade lidars, lidars on consumer devices have lower spatial and temporal resolution. We highlight an advantage of PlatoNeRF over BF Lidar, which is better generalization to low-resolution regimes due to our implicit representation. Quantitative and qualitative results are shown in Tab. 3 and Fig. 6 (rows 1–2), respectively.
@@ 4.3.2 Ambient Light and Low Albedo Backgrounds
In real-world settings, there may be high ambient light or low scene albedo, both of which make detection of shadows in RGB images challenging. In contrast, lidar-based methods, such as PlatoNeRF, are fundamentally more robust to these low signal-to-noise (SNR) and signal-to-background (SBR) scenarios. While ambient light and low albedo impact both RGB and lidar-based approaches, lidar-based approaches can mitigate their impact with time- and wavelength-gating. Ambient noise is uniformly distributed over time. Time gating enables suppression of ambient noise by only considering timing bins containing the pulse signal. Wavelength gating enables suppression of ambient noise in wavelengths not measured by the SPAD sensor, whereas RGB sensors have broadband sensitivity to visible wavelengths. We provide a detailed explanation of the gating principle in the supp. Quantitative results for these ablations are reported in Tab. 4.
@@
Increasing ambient light increases the background and noise in the SBR or SNR term. We empirically validate that PlatoNeRF is able to handle ambient light in the scene, while $\text{S}^{3}$-NeRF depth error increases. For this experiment, we render the scene with an added area light to train both methods. The area light intensity is the same in both RGB and lidar rendering. We do not model saturation or pileup distortion [30] effects, since these are not significant indoors. PlatoNeRF’s reconstruction under ambient light is shown in Fig. 6 (row 3) and $\text{S}^{3}$-NeRF’s is in the supp.
@@
Reducing albedo reduces the signal in the SNR and SBR term. However, unlike RGB sensors, albedo has a minor effect on lidar. To show its impact on PlatoNeRF and $\text{S}^{3}$-NeRF, we reduce the albedo of all background surfaces. While the shadow is still discernible to the human eye (see supp.), the lowered albedo causes $\text{S}^{3}$-NeRF’s depth error to increase, while PlatoNeRF is unaffected. $\text{S}^{3}$-NeRF produces accurate depth from the training view, but geometry in occluded regions is missing due to the weaker shadows.
@@ 4.3.3 Other Ablations
Lastly, we study the impact of (a) non-planar background geometry, (b) number of illumination spots, and (c) shadow mask threshold on PlatoNeRF, with additional results for each in the supp. Qualitative results for (a) and (b) are shown in Fig. 6 (row 3). The non-planar scene contains curved background walls. PlatoNeRF’s depth L1 is 7.75 cm and PSNR is 27.25 dB across test views, similar to previous experiments. For our illumination spot ablation, we reduce the number of illumination spots from 16 to 8, leading to L1 error of 9.12 cm and PSNR of 26.33 dB across test views, a small drop from when all 16 views are used. For our shadow mask threshold ablation, we vary the threshold between zero and one, resulting in L1 errors between 6.88 and 70 cm, highlighting the importance of the shadow mask threshold.
@@ 5 Conclusion
We believe this research is a promising direction as lidar becomes ubiquitous. Future directions enabled by PlatoNeRF include incorporating RGB and lidar with neural rendering to recover textured geometry, using data priors often used in single-view 3D reconstruction [49], handling specularities that cause ambiguities in the time of flight, and modeling more than two-bounces of light.
@@ Appendix A Time & Wavelength Gating in Lidar
The principle of time gating is illustrated in Fig. 7. A measured lidar signal $i(t)$ can be decomposed into the pulse signal $s(t)$ and (roughly) constant ambient background noise $n(t)=N$. An RGB sensor would integrate over this timing information and measure
@@
where $T_{1}$ and $T_{2}$ determine the gated window in the transient signal and $W=T_{2}-T_{1}$ is the window size. Note that the numerator of Eq. 11 is roughly the same as the SBR in Eq. 13 because $s(t)\approx 0$ for $t<T_{1}$ and $t>T_{2}$, as shown in Fig. 7(a). Therefore, time gating offers an SNR improvement of $\frac{T}{W}$ over techniques that leverage RGB or intensity signals. Note that the SBR enhancement is inversely proportional to the gated window. We do not account for Poisson noise effects, which, in practice, would introduce trade-offs in determining the window size. Empirical results are plotted in Fig. 7(b)-(d) on the effects of time gating on enhancing contrast in shadow images.
@@
A similar idea can be applied to gate wavelengths. Most of the signal will be concentrated within a narrow spectral range, and all other intensities can be gated out with a narrow-band pass filter, as shown in Fig. 8. This figure plots the emission spectra of an LED light [36] and the gating profile is determined by a $685$ nm PicoQuant pulsed laser [31].
@@ Appendix B Simulated Dataset Details
In this section, we describe the simulated datasets that we render and use to compare our method to past work in more detail. We render four simulated scenes, as described in the main text, with both a lidar and RGB camera in Mitsuba [12]. The lidar data is used to run PlatoNeRF and Bounce Flash (BF) Lidar [7] and the RGB data is used to run $\text{S}^{3}$-NeRF [47]. The same sixteen scene points are illuminated in both the lidar and RGB data. In the lidar data, the sixteen points are illuminated with a laser and, in the RGB data, point light sources are placed at each of the sixteen points. A camera to world transform from OpenGL (x right, y up, z back) to Mitsuba (x left, y up, and z forward) is used to train each method with this data. Ground truth depth for both the train view and 120 test views are provided. A subset of the test view frames are shown in the video results on the project page. All data has been released for use in future work.
@@ Appendix C Training Details
We have released all data, code, and model checkpoints, along with documentation, to ensure our work is fully reproducible by others. These can be accessed from our project webpage. Simulated data is rendered in Mitsuba world coordinates and PlatoNeRF uses OpenGL camera coordinates. Code for simulating time of flight measurements in Mitsuba is also provided.
@@
We train our model for 200k iterations. For the first 25k iterations, only the distance loss is applied, while both the distance and shadow losses are applied thereafter. We use a threshold of 15% on the shadow confidence map (computed as the maximum of the cross-correlation described in Sec 3.3 of the main text) when extracting ground truth shadow masks from the raw lidar measurements. This threshold is used across all experiments, except the ambient light experiment, where we further tune it.
@@ Appendix D Extended Ablations
In this section, we add further detail and discussion on the results of our ablations, quantitatively reported in the main text. In addition, we provide further ablation on the impact of non-planar background geometry (Fig. 11), the number of illumination points (Tab. 5), and the shadow mask threshold (Fig. 13) on PlatoNeRF reconstruction.
@@ D.1 Spatial- and Temporal-Resolution
Figure 10: Ambient Light and Low Albedo Background Ablation. We compare PlatoNeRF and $\text{S}^{3}$-NeRF when trained on scenes with ambient light or low albedo background. PlatoNeRF is robust to both, whereas the performance of $\text{S}^{3}$-NeRF degrades.
Qualitative results comparing PlatoNeRF and Bounce Flash (BF) Lidar under varying spatial- and temporal-resolutions are shown in Fig. 9. This ablation is important because lidars on consumer devices are often constrained to much lower resolutions than research-grade lidars. Spatial resolution is varied by downsampling the number of pixels, while keeping the field of view of the lidar the same. As spatial resolution is reduced, geometry predicted by BF Lidar becomes sparser. The depth estimation of visible points in the scene remains accurate, but there is no interpolation between these points. The sparsity in visible depth information negatively impacts the shadow carving step of BF Lidar, leading to poor reconstruction of the chair in lower spatial resolution regimes. On the other hand, because PlatoNeRF is able to smoothly interpolate across missing pixels, the resulting reconstruction is significantly more accurate.
@@ D.2 Ambient Light
Qualitative results comparing PlatoNeRF and $\text{S}^{3}$-NeRF reconstructions under ambient light are shown in Fig. 10 (top row). While $\text{S}^{3}$-NeRF is able to model small amounts of ambient light, it fails under realistic amounts of ambient light, in this case, from an added area light. On the other hand, PlatoNeRF is still able to accurately reconstruct the scene with the same ambient light added.
@@ D.3 Low-Albedo Backgrounds
Qualitative results comparing PlatoNeRF and $\text{S}^{3}$-NeRF reconstructions with a low albedo background are shown in Fig. 10 (bottom row). $\text{S}^{3}$-NeRF is able to accurately reconstruct the visible portion of the scene, but is unable to recover occluded geometry due to worse contrast in the shadow (though it is still discernible to the human eye, as shown in Fig. 10). On the other hand, PlatoNeRF is not significantly affected by scene albedo due to its use of a lidar rather than RGB sensor.
Figure 11: Non-Planar Background Geometry Ablation. We provide an additional example of PlatoNeRF reconstruction of a scene with complex, non-planar background geometry. PlatoNeRF accurately reconstructs both the non-planar foreground and background geometery from a single view.
Figure 13: Shadow Mask Ablation. We vary the threshold used when creating shadow masks and report the change in L1 depth error across all test viewpoints. Ablation is done on the chair scene.
D.4 Non-Planar Background Geometry
We study the impact of non-planar background geometry on PlatoNeRF. In the main text, we show results on a scene with curved walls, resulting in similar depth L1 and PSNR scores as the same scene with planar walls. This result indicates that PlatoNeRF is robust to non-planar foreground and background geometry. In Fig. 11, we further increase the complexity of the background geometry by adding two objects: a couch and painting. As shown by the extracted shadow in Fig. 11, the additional background objects cause the shadow to be contorted based on the geometry its cast on. PlatoNeRF is still able to accurately reconstruct the full scene geometry. Certain parts of the scene, such as the wall behind the chair and the self-occluded face of the armrest, will not have two-bounce ToF measurements, resulting in PlatoNeRF interpolating the geometry in those areas.
D.5 Number of Illumination Points
We further ablate the impact of reducing the number of illumination points used to train PlatoNeRF. In our main experiments, we use sixteen illumination points. We reduce that number to eight, four, and two and report the results in Tab. 5. Qualitative results are shown in Fig. 12. The scene is reconstructed for each number of illumination points, however, as the number is reduced, quality also decreases, as there is less information about occluded regions. When there are only two illumination points, the occluded chair legs are not reconstructed. We note that while we study the number of illumination points, their location is also an important factor in reconstruction quality. As the number of illumination points is reduced, the location of the remaining illumination points becomes increasingly important, i.e. casting shadows with the most relevance and diversity. In these experiments, we randomly choose which illumination points to use.
D.6 Shadow Mask Threshold
We ablate the impact of the shadow mask threshold on PlatoNeRF reconstruction quality. Shadow masks are generated from the raw time-of-flight data, as described in Sec 3.3 of the main text. To ablate the impact of the probability threshold used to extract shadow masks, we vary it between zero and one in increments of $0.125$. Fig. 13 shows the resulting shadow masks and depth L1 error across test views at each threshold. While a shadow probability threshold of $0.15$ was used to generate the results in the main text, ablation results indicate that a threshold of $0.25$ leads to even better performance. While the approach employed by PlatoNeRF for extracting shadow masks is common in past work, such as BF Lidar, PlatoNeRF is agnostic to the shadow segmentation approach and more sophisticated methods [41] can be extended to transient data and employed in the future.